# Optimal multiple-phase estimation with multi-mode NOON states against photon loss

### v1

@@ Appendix A. Phase-encoded multi-mode NOON states under photon loss
with the vacuum state $|0\rangle$ in the environment mode $j$ and unitary operator $\hat{B}_{j}$ with respect to a mode transformation $\hat{B}_{j}\hat{a}_{j}^{\dagger}\hat{B}_{j}^{\dagger}=\sqrt{1-\gamma_{j}}\hat{a}_{j}^{\dagger}+\sqrt{\gamma_{j}}\hat{e}_{j}^{\dagger}$ s.m.barnett , where $\mathrm{Tr}_{e_{j}}\left\{\cdot\right\}$ is the partial trace over the $j$-th environment system. From Eq. (20), It is straightforward to derive three terms $\Lambda_{\gamma_{j}}(|N\rangle\langle N|)$, $\Lambda_{\gamma_{j}}(|N\rangle\langle 0|)$, and $\Lambda_{\gamma_{j}}(|0\rangle\langle 0|)$ as
@@
$\hat{\rho}_{N,\bm{\phi}}=\left\{\sum_{j=0}^{d}p_{j}(1-\gamma_{j})^{N}\right\}|\widetilde{\psi}_{N,\bm{\phi}}\rangle\langle\widetilde{\psi}_{N,\bm{\phi}}|+\underbrace{\sum_{j=0}^{d}p_{j}\left\{\sum_{r=1}^{N}\frac{N!}{r!(N-r)!}(1-\gamma_{j})^{N-r}\gamma_{j}^{r}\frac{\hat{a}_{j}^{\dagger N-r}|0\rangle\langle 0|^{\otimes d}\hat{a}_{j}^{N-r}}{(N-r)!}\right\}}_{\hat{\sigma}_{N}}.$ (22)
@@ Appendix B. Evaluation of the minimum QCRB
Here we analytically evaluate the minimum QCRB of $d+1$-mode NOON states. The non-negative operator $\hat{\sigma}_{N}$ in Eq. (22) is diagonalized by vectors of the Fock basis whose photon number is strictly less than $N$. Meanwhile, $|\widetilde{\psi}_{N,\bm{\phi}}\rangle$ in Eq. (22) is a linear combination of the Fock basis vectors with photon number equal to $N$. Thus, the support of $\hat{\sigma}_{N}$ in Eq. (22) is orthogonal to $|\widetilde{\psi}_{N,\bm{\phi}}\rangle$, and $\hat{\rho}_{N,\bm{\phi}}$ is in the spectral decomposition with the eigenvectors $|0\rangle,\cdots,|N-1\rangle$ together with $|\widetilde{\psi}_{N,\bm{\phi}}\rangle$. We can employ the theoretical framework written in Ref. j.liu2 to derive components of the QFIM
@@
which is the convex function of $\bm{p}=(p_{0},\cdots,p_{d})$. Since a feasible set of $\bm{p}$ is convex, a global minimum is at the point where gradient of the QCRB in Eq. (24) is vanished s.boyd . Hence, the minimum QCRB is evaluated as Eq. (10) by the optimal weights in Eq. (III.1).
@@ Appendix C. Attainability of the QCRB
where $\hat{L}_{a}$ and $\hat{L}_{b}$ are symmetric logarithm derivative operators. By using the above proposition, we show that there is a measurement attaining the QCRB in Eq. (24). From the output state described by Eq. (22), the trace of the left hand side in Eq. (25) is written as linear combination of $\langle\widetilde{\psi}_{N,\bm{\phi}}|[\hat{L}_{a},\hat{L}_{b}]|\widetilde{\psi}_{N,\bm{\phi}}\rangle$ and $\langle 0|^{\otimes d+1}\hat{a}_{j}^{N-r}[\hat{L}_{a},\hat{L}_{b}]\hat{a}_{j}^{\dagger N-r}|0\rangle^{\otimes d+1}$ with $r\not=0$. First, since $|\widetilde{\psi}_{N,\bm{\phi}}\rangle$ in Eq. (22) is a multi-mode NOON state, there are symmetric logarithm derivative operators $\hat{L}_{a}$ and $\hat{L}_{b}$ such that $\langle\widetilde{\psi}_{N,\bm{\phi}}|[\hat{L}_{a},\hat{L}_{b}]|\widetilde{\psi}_{N,\bm{\phi}}\rangle=0$ p.c.humphreys . Second, as we consider $\hat{L}_{a}$ and $\hat{L}_{b}$ of the multi-mode NOON state $|\widetilde{\psi}_{N,{\bm{\phi}}}\rangle$, their supports are orthogonal to $\hat{a}_{j}^{\dagger N-r}|0\rangle^{\otimes d+1}$ as $r\not=0$. Thus, $\langle 0|^{\otimes d+1}\hat{a}_{j}^{N-r}[\hat{L}_{a},\hat{L}_{b}]\hat{a}_{j}^{\dagger N-r}|0\rangle^{\otimes d+1}$ also become zero. Finally, the symmetric logarithm derivative operators satisfy the equality in Eq. (25), which leads us to understand the attainability of the QCRB.
@@ Appendix D. Evaluation of the SQL
Here, we evaluate the SQL quantified by the QCRB of multi-mode coherent states. If the mean photon number of an initial coherent states $\bigotimes_{j=0}^{d}|\alpha_{j}\rangle$ with $\alpha_{j}=\sqrt{q_{j}N}$ is not sufficiently large, then the coherent state in the reference mode is not understood as strong. Thus, it is reasonable to describe a probe state as a phase-averaged multi-mode coherent state a.z.goldberg ; m.jarzyna :
[…]
with $\beta_{j}=\alpha_{j}e^{i\phi_{j}}$. Of course, the above probe state encapsulates the case of the strong reference beam when $|\alpha|^{2}$ is sufficiently large and $|\beta_{0}|^{2}$ is much larger than any $|\beta_{j}|^{2}$ with $j\not=0$ m.jarzyna . Photon loss channels transform the above phase-averaged multi-mode coherent state to
[…]
Finally, the QCRB of the above noisy state is analytically evaluated as a.z.goldberg
$\mathrm{Tr}\left\{{F}_{\rm Q,coh}^{-1}(\bm{\phi})\right\}=\frac{d}{4(1-\gamma_{0})|\beta_{0}|^{2}}+\frac{1}{4}\sum_{j=1}^{d}\frac{1}{(1-\gamma_{j})|\beta_{j}|^{2}}=\frac{1}{4N}\left\{\frac{d}{(1-\gamma_{0})q_{0}}+\sum_{j=1}^{d}\frac{1}{(1-\gamma_{j})q_{j}}\right\}.$ (28)
Thus, the minimum value of the QCRB over all $q_{j}$ satisfying the relation $\sum_{j=0}^{d}q_{j}=1$ is straightforwardly evaluated as Eq. (10).

### v2

@@ Appendix A. Phase-encoded multi-mode NOON states under photon loss
with the vacuum state $|0\rangle$ in the environment mode $j$ and unitary operator $\hat{B}_{j}$ with respect to a mode transformation $\hat{B}_{j}\hat{a}_{j}^{\dagger}\hat{B}_{j}^{\dagger}=\sqrt{1-\gamma_{j}}\hat{a}_{j}^{\dagger}+\sqrt{\gamma_{j}}\hat{e}_{j}^{\dagger}$ s.m.barnett , where $\mathrm{Tr}_{e_{j}}\left\{\cdot\right\}$ is the partial trace over the $j$-th environment system.
From Eq. (22), It is straightforward to derive three terms $\Lambda_{\gamma_{j}}(|N\rangle\langle N|)$, $\Lambda_{\gamma_{j}}(|N\rangle\langle 0|)$, and $\Lambda_{\gamma_{j}}(|0\rangle\langle 0|)$ as
@@
$\hat{\rho}_{N,\bm{\phi}}$ =$=\left\{\sum_{j=0}^{d}p_{j}(1-\gamma_{j})^{N}\right\}|{\psi}_{N,\bm{\phi}}^{(\bm{\gamma})}\rangle\langle{\psi}_{N,\bm{\phi}}^{(\bm{\gamma})}|$
$+\underbrace{\sum_{j=0}^{d}p_{j}\left\{\sum_{r=1}^{N}\frac{N!}{r!(N-r)!}(1-\gamma_{j})^{N-r}\gamma_{j}^{r}\frac{\hat{a}_{j}^{\dagger N-r}|0\rangle\langle 0|^{\otimes d}\hat{a}_{j}^{N-r}}{(N-r)!}\right\}}_{\hat{\sigma}_{N}}.$ (24)
@@ Appendix B. Evaluation of the minimum QCRB
We begin the detailed evaluation by introducing methodology to evaluate the QFIM, which was proposed in Refs. j.liu2 ; m.g.a.paris ; j.liu_sld . Let us consider a density operator $\hat{\rho}_{\bm{\phi}}=\sum_{j}\lambda_{j}|\lambda_{j}\rangle\langle\lambda_{j}|$ with eigenvalues $\lambda_{j}$ and corresponding eigenvectors $|\lambda_{j}\rangle$. If $\hat{\rho}_{\bm{\phi}}$ is full rank, then an SLD operator defined in Eq. (5) is decomposed as m.g.a.paris
$\hat{L}_{a}=\sum_{j}\frac{\partial_{\phi_{a}}\lambda_{j}}{\lambda_{j}}|\lambda_{j}\rangle\langle\lambda_{j}|+2\sum_{j\not=k}\frac{\lambda_{j}-\lambda_{k}}{\lambda_{j}+\lambda_{k}}\langle\lambda_{k}|\partial_{\phi_{a}}\lambda_{j}\rangle|\lambda_{k}\rangle\langle\lambda_{j}|.$ (25)
The above decomposition admits Eq. (5) to be written as j.liu2 ; j.liu_sld
$F_{\mathrm{Q}(a,b)}(\bm{\phi})$ $=$ $\sum_{j}\frac{(\partial_{\phi_{a}}\lambda_{j})(\partial_{\phi_{b}}\lambda_{j})}{\lambda_{j}}+4\sum_{j}\lambda_{j}\mathrm{Re}\left\{\langle\partial_{\phi_{a}}\lambda_{j}|\partial_{\phi_{b}}\lambda_{j}\rangle\right\}$ (26)
-- $8\sum_{j,k}\frac{\lambda_{j}\lambda_{k}}{\lambda_{j}+\lambda_{k}}\mathrm{Re}\left\{\langle\partial_{\phi_{a}}\lambda_{j}|\lambda_{k}\rangle\langle\lambda_{k}|\partial_{\phi_{b}}\lambda_{j}\rangle\right\}.$
Based on the QFIM formalism introduced above, we analytically evaluate the minimum QCRB of $d+1$-mode NOON states. The non-negative operator $\hat{\sigma}_{N}$ in Eq. (Appendix A. Phase-encoded multi-mode NOON states under photon loss) is diagonalized by vectors of the Fock basis whose photon number is strictly less than $N$. Meanwhile, $|{\psi}_{N,\bm{\phi}}^{(\bm{\gamma})}\rangle$ in Eq. (Appendix A. Phase-encoded multi-mode NOON states under photon loss) is a linear combination of the Fock basis vectors with photon number equal to $N$. Thus, the support of $\hat{\sigma}_{N}$ in Eq. (Appendix A. Phase-encoded multi-mode NOON states under photon loss) is orthogonal to $|{\psi}_{N,\bm{\phi}}^{(\bm{\gamma})}\rangle$, and $\hat{\rho}_{N,\bm{\phi}}$ is in the spectral decomposition with the eigenvectors $|0\rangle,\cdots,|N-1\rangle$ together with $|{\psi}_{N,\bm{\phi}}^{(\bm{\gamma})}\rangle$. We can obtain Eq. (III.1) from Eq. (26), which is employed to derive components of the QFIM
@@
which is the convex function of $\bm{p}=(p_{0},\cdots,p_{d})$. Since a feasible set of $\bm{p}$ is convex and $\mathrm{Tr}\left\{F_{\rm Q}^{-1}(\bm{\phi})\right\}$ is a convex function, a global minimum is represented as a point where gradient of the QCRB in Eq. (28) is vanished s.boyd . Also, normalization condition $p_{0}+p_{1}+\cdots+p_{d}=1$ leads us to
$\frac{\partial}{\partial p_{j}}\mathrm{Tr}\left\{F_{\rm Q}(\bm{\phi})\right\}=\frac{1}{4N^{2}}\left\{\frac{d}{p_{0}^{2}(1-\gamma_{0})^{N}}-\frac{1}{p_{j}^{2}(1-\gamma_{j})^{N}}\right\}$ (29)
with vanishing-gradient conditions
$p_{j}=\frac{1}{\sqrt{d}}\frac{(1-\gamma_{0})^{N/2}}{(1-\gamma_{j})^{N/2}}p_{0},\ \ \forall j\in\{1,\cdots,d\}.$ (30)
Hence, we obtain optimal weights in Eq. (III.1) from $p_{j}$ written above together with the normalization condition. Consequently, the minimum QCRB is evaluated as Eq. (12) by the optimal weights.
@@ Appendix C. Attainability of the QCRB
where $\hat{L}_{a}$ and $\hat{L}_{b}$ are symmetric logarithm derivative operators. By using the above proposition, we show that there is a measurement attaining the QCRB in Eq. (28). From the output state described by Eq. (Appendix A. Phase-encoded multi-mode NOON states under photon loss), the trace of the left hand side in Eq. (31) is written as linear combination of $\langle{\psi}_{N,\bm{\phi}}^{(\bm{\gamma})}|[\hat{L}_{a},\hat{L}_{b}]|{\psi}_{N,\bm{\phi}}^{(\bm{\gamma})}\rangle$ and $\langle 0|^{\otimes d+1}\hat{a}_{j}^{N-r}[\hat{L}_{a},\hat{L}_{b}]\hat{a}_{j}^{\dagger N-r}|0\rangle^{\otimes d+1}$ with $r\not=0$. First, since $|{\psi}_{N,\bm{\phi}}^{(\bm{\gamma})}\rangle$ in Eq. (Appendix A. Phase-encoded multi-mode NOON states under photon loss) is a multi-mode NOON state, there are symmetric logarithm derivative operators $\hat{L}_{a}$ and $\hat{L}_{b}$ such that $\langle{\psi}_{N,\bm{\phi}}^{(\bm{\gamma})}|[\hat{L}_{a},\hat{L}_{b}]|{\psi}_{N,\bm{\phi}}^{(\bm{\gamma})}\rangle=0$ p.c.humphreys . Second, as we consider $\hat{L}_{a}$ and $\hat{L}_{b}$ of the multi-mode NOON state $|{\psi}_{N,\bm{\phi}}^{(\bm{\gamma})}\rangle$, their supports are orthogonal to $\hat{a}_{j}^{\dagger N-r}|0\rangle^{\otimes d+1}$ as $r\not=0$. Thus, $\langle 0|^{\otimes d+1}\hat{a}_{j}^{N-r}[\hat{L}_{a},\hat{L}_{b}]\hat{a}_{j}^{\dagger N-r}|0\rangle^{\otimes d+1}$ also become zero. Finally, the symmetric logarithm derivative operators satisfy the equality in Eq. (31), which leads us to understand the attainability of the QCRB.
@@ Appendix D. Evaluation of the SQL
We first remind that total uncertainty $(\Delta\bm{\phi})^{2}$ is equal to $(\Delta\bm{\phi})^{2}=\sum_{j=1}^{d}(\Delta(\widetilde{\phi}_{j}-\widetilde{\phi}_{0}))^{2}=d(\Delta\widetilde{\phi}_{0})^{2}+\sum_{j=1}^{d}(\Delta\widetilde{\phi}_{j})^{2}$, and each $(\Delta\widetilde{\phi}_{j})^{2}$ with respect to a lossy coherent state $|e^{i\widetilde{\phi}_{j}}\sqrt{1-\gamma_{j}}\alpha_{j}\rangle$ admits QCRB inequality $(\Delta\widetilde{\phi}_{j})^{2}\geq\frac{1}{4|\alpha_{j}|^{2}(1-\gamma_{j})}$ c.oh . Thus, lower bound of the total uncertainty is evaluated as
$(\Delta\bm{\phi})^{2}$ $=$ $d(\Delta\widetilde{\phi}_{0})^{2}+\sum_{j=1}^{d}(\Delta\widetilde{\phi}_{j})^{2}$ (32)
$\geq$ $\frac{d}{4|\alpha_{0}|^{2}(1-\gamma_{0})}+\frac{1}{4|\alpha_{1}|^{2}(1-\gamma_{1})}+\cdots+\frac{1}{4|\alpha_{d}|^{2}(1-\gamma_{d})}$
$=$ $\frac{1}{4N}\left\{\frac{d}{q_{0}(1-\gamma_{0})}+\frac{1}{q_{1}(1-\gamma_{1})}+\cdots+\frac{1}{q_{d}(1-\gamma_{d})}\right\},$
where we parameterize $|\alpha_{j}|$ as $|\alpha_{j}|=\sqrt{q_{j}N}$ with $q_{0}+\cdots+q_{d}=1$ and $q_{j}\geq 0$ $\forall j\in\{0,\cdots,d\}$. It is straightforward to show that Hessian matrix of the lower bound in Eq. (32) is positive-definite for any points $(q_{0},\cdots,q_{d})$, which means that a point that vanishes the gradient yields the global minimum s.boyd . Consequently, the minimum QCRB written in Eq. (12) is obtained by substituting the vanishing-gradient point to the lower bound in Eq. (32).
@@ Appendix E. Measurement probabilities for the CRB
Figure 7: Histogram of multi-mode NOON states’ QCRB in cases with arbitrarily chosen photon loss rates $\gamma_{j}\in[0.2,0.6]$.
Appendix F. Investigation of QCRB in cases with general photon loss rates
For considering general loss cases, we consider $10^{5}$ instances with arbitrarily chosen photon loss rates from a uniform distribution within a specific range $[\gamma_{\rm min}=0.2,\gamma_{\rm max}=0.6]$, particularly for cases where $N=2$ and $d=10$. For individual instances, we calculate the QCRBs for our optimal NOON states, Humphreys et al.’s NOON states p.c.humphreys , and coherent states, and illustrate the results in Fig. 7. We observe that the distribution of the optimal multi-mode NOON states is more left-skewed than that of the multi-mode NOON states from Ref. p.c.humphreys , although both distributions have the common minimum and maximum when $\gamma_{j}=\gamma_{\rm min}$ and $\gamma_{j}=\gamma_{\rm max}$ for all $j$’s, respectively. It means that estimation schemes are more robust to photon loss in general since our NOON states are optimized individually with respect to photon loss rates. We also note that the distribution of NOON states (both ours and Humphreys’ et al.’s p.c.humphreys ) is wider than that of multi-mode pure coherent states. This implies that multi-mode NOON states are more sensitive against photon loss than the coherent states.